OpenStudy (anonymous):

Jennifer has been saving for college for 57 months. The first month, she saved $11. She was able to save more money each month than the month before. She ended up saving $19,779.00. How much more did she save each month?

OpenStudy (tkhunny):

First, there is not enough information. The problem cannot be solved as stated. Second, let's assume the increase in savings was the same each month. Month 1: 11 Month 2: 11+m Month 3: 11+2m Month 4: 11+3m ... Month 57: 11 + 56m Can you add those up? There might be an Arithmetic Series in there, somewhere.

OpenStudy (anonymous):

How do you want me to do it is 11+56m=67

OpenStudy (tkhunny):

?? Why would m = 1? You must add up ALL the months. 11 + 11+m + 11+2m + 11+3m + ... + 11 + 56m = 19779 57(11) + (m + 2m + 3m + ... + 56m) = 19779 m(1 + 2 + 3 + ... + 56) = 19779 - 57(11) = 19779 - 627 = 19152 m(56)(57)/2 = 19779 - 57(11) = 19779 - 627 = 19152 m = 19152*2/(56*57) -- I hope this comes out to be an integer! m = 12 -- Woo hoo!! 11 + 23 + 35 + ... + 683 = 19779!!

OpenStudy (anonymous):

Thank you so much i did it i found 12 for it.. Thank you again for the help